# Reconstruction of Caterpillar Tanglegrams

### v1

@@ 4. Reconstruction of caterpillar tanglegrams
Lemmas 4.3, 4.4, and 4.5 establish our main theorem, which we re-state for the reader’s convenience.
@@
We end this manuscript by noting that Theorem 1.1 establishes the reconstruction of caterpillar tanglegrams of size $n\geq 5$ from the multideck $\mathcal{D}^{m}(\mathcal{T})$. Indeed, our proofs rely on the multiplicities of cards in several instances. However, since trees of size $n\geq 6$ are reconstructable from their decks (Theorem 2.7), it is an interesting open question to investigate whether decks also determine sufficiently large tanglegrams (note that we must have $n\geq 6$ for the trees to be reconstructable from the deck).
Problem 4.6.
Determine whether sufficiently large tanglegrams are reconstructable from their decks.
@@ Acknowledgements
This project was initiated when the authors participated in the Mathematics Research Community workshop “Trees in Many Contexts” (Beaver Hollow, NY), organized by the American Mathematical Society with the support of the National Science Foundation under Grant Number DMS-1641020. The work on this project was continued when EC, LS, and KW were in residence at the Institute for Computational and Experimental Research in Mathematics of Brown University (Providence, RI), during the “Theory, Methods, and Applications of Quantitative Phylogenomics” semester program supported by the National Science Foundation under Grant Number DMS-1929284.

### v2

@@ 8. Conclusion
Symmetry arguments and Lemmas 5.3, 6.3, and 7.3 establish our main theorem, which we re-state for the reader’s convenience.
@@
We note that Theorem 1.1 establishes the reconstruction of caterpillar tanglegrams of size $n\geq 6$ from the multideck $\mathcal{D}^{m}(\mathcal{T})$. Indeed, our proofs rely on the multiplicities of cards in several instances. However, since trees of size $n\geq 6$ are reconstructable from their decks (Lemma 2.17), it is an interesting open question to investigate whether decks also determine sufficiently large tanglegrams of some tanglegram class.
The original question — whether tanglegrams are reconstructable for sufficiently large size — remains open. Interestingly, not all tanglegrams of size $6$ are reconstructable. An example of a pair of tanglegrams of size $6$ that have the same multideck is shown in Figure 8. Using computer search [14], we found $4$ pairs of size-$6$ tanglegrams that have the same multideck. In each of these tanglegrams, the left and right trees were a $C_{3}\oplus C_{3}$, and either $C_{2}\oplus(C_{2}\oplus C_{2})$ or $(C_{1}\oplus)^{2}(C_{2}\oplus C_{2})$.
$v_{1}$
$v_{2}$
$v_{3}$
$u_{4}$
$v_{5}$
$v_{6}$
$u_{1}$
$u_{2}$
$u_{3}$
$u_{4}$
$u_{5}$
$u_{6}$
$\mathcal{T}_{1}$
$v_{1}$
$v_{2}$
$u_{3}$
$u_{4}$
$v_{5}$
$v_{6}$
$u_{1}$
$u_{2}$
$u_{3}$
$u_{4}$
$u_{5}$
$u_{6}$
$\mathcal{T}_{2}$
$\mathcal{T}_{1}-v_{2}u_{1}=\mathcal{T}_{1}-v_{5}u_{2}$$\mathcal{T}_{2}-v_{2}u_{1}=\mathcal{T}_{2}-v_{5}u_{2}$$\mathcal{T}_{1}-v_{1}u_{3}$$\mathcal{T}_{2}-v_{3}u_{3}$$\mathcal{T}_{1}-v_{3}u_{5}$$\mathcal{T}_{2}-v_{1}u_{5}$
$\mathcal{T}_{1}-v_{4}u_{4}$$\mathcal{T}_{2}-v_{6}u_{6}$$\mathcal{T}_{1}-v_{6}u_{6}$$\mathcal{T}_{2}-v_{4}u_{4}$
Figure 8. Two non-isomorphic size $6$ tanglegrams that have the same multideck, and the cards in their deck. A labeling of the left and right trees of size $6$ tanglegrams are given as well as all representations of the cards in the deck.
In general, we ask the following:
Problem 8.1.
Given a tanglegram class $\mathcal{H}$, determine whether sufficiently large tanglegrams in the class are reconstructable from their deck or multideck. If yes, what is the best threshold for “sufficiently large”?
The requirement of reconstructibility here includes being able to identify membership in $\mathcal{H}$ from the deck or multideck (see Definition 2.15). Theorem 1.1 is of this type: $\mathcal{H}$ is the set of tanglegrams where at least one of the left and right tree is a caterpillar. In this case (and in any case when membership in $\mathcal{H}$ is determined by identifying left and right trees) being able to identify the membership relation from the multideck or the deck is not a significant requirement (see Lemma 2.17).
A class of tanglegrams where left and right trees alone do not determine membership in the class is the class of planar tanglegrams, that is, tanglegrams that have a layout without any crossing edges. Czabarka, Wagner and Székely [6] showed a tanglegram analogy of Kuratowski’s theorem: every nonplanar tanglegram contains one of two size $4$ subtanglegrams. This result implies that given a deck of any tanglegram with size different from $4$, one can identify from its deck whether it is planar or not. Interestingly, none of the tanglegrams among the size $6$ tanglegrams that are not reconstructable from their multideck are planar, therefore it is possible that all size at least $6$ planar tanglegrams are reconstructable from their multideck. The authors are preparing a manuscript on the reconstructibility of planar tanglegrams.
@@ Acknowledgements
This project was initiated when the authors participated in the Mathematics Research Community workshop “Trees in Many Contexts” (Beaver Hollow, NY), organized by the American Mathematical Society with the support of the National Science Foundation under Grant Number DMS-1641020. The work on this project was continued when EC, LS, and KW were in residence at the Institute for Computational and Experimental Research in Mathematics of Brown University (Providence, RI), during the “Theory, Methods, and Applications of Quantitative Phylogenomics” semester program supported by the National Science Foundation under Grant Number DMS-1929284. We thank two anonymous reviewers for their valuable feedback on an earlier version of this manuscript.